I'm trying to help my daughter with this problem and we both come up with different answers.
-6x+3(4x-1)=9 What is X

Answer by askmemath(368)

-6x+3(4x-1)=9 What is X
In order to remove the brackets, we multiply every term within the brackets with 3.So we get
-6x + 12x -3 = 9
Adding +3 on both sides
6x= 9+3
6x = 12
Dividing by 6 throughout
x = 2